OpenStudy (anonymous):

Multiply and simplify: x+11/x+3 x^2-9/ x^2+20x+99

OpenStudy (paxpolaris):

\[\Large{ x+11\over x+3}\times{ x^2-9\over x^2+20x+99}\]???

OpenStudy (anonymous):

yess

OpenStudy (paxpolaris):

factor the numerator & denom. on the right

OpenStudy (paxpolaris):

\[\Large={ x+11\over x+3}\times{ \left( x+3 \right)\left( x-3 \right)\over \left( x+11 \right)\left( x+9 \right)}\]

OpenStudy (paxpolaris):

\[\Large={ \cancel{x+11}\over \cancel{x+3}}\times{ \cancel{\left( x+3 \right)}\left( x-3 \right)\over \cancel{\left( x+11 \right)}\left( x+9 \right)}\]

OpenStudy (paxpolaris):

\[\huge={x-3 \over x+9}\]